OpenStudy (anonymous):

Multiply. Write the product in simplest form. √2(√8+√4)

OpenStudy (anonymous):

\[\sqrt{2}(2\sqrt{2}+2)\rightarrow 2(\sqrt{2})^2+2\sqrt{2}\rightarrow 4+2\sqrt{2}\]

OpenStudy (anonymous):

since \[\sqrt{8}=\sqrt{4*2}=\sqrt{2^2*2}=2\sqrt{2}\]

OpenStudy (anonymous):

@smith1512 knock wrote the answer above is there a step you do not understand?

OpenStudy (anonymous):

\[\sqrt{2}(\sqrt{8}+\sqrt{4})=\sqrt{16}+\sqrt{8}=4+2\sqrt{2}\]

OpenStudy (anonymous):

Thanks a bunch guys <3 means alot